OpenStudy (newwar):

Which ordered pairs are on the line with equation ​ y−0.75=1.25(x−3)? Select each correct answer. ​ (−4, −8) ​ ​ (7, 5.5) ​ ​ (1, 1.75) ​ ​ (4, 2) ​

OpenStudy (newwar):

i think it is C and D

OpenStudy (angel.candy):

The answer is Option D. (4,2) in this type of question , you just need to input the (x,y) , and see if the amount is correct y - 0.75 = 1.25 (x-3) 2 - 0.75 = 1.25 ( 4 - 3) 1.25 = 1.25 (1) 1.25 = 1.25